In addition to these immediate issues, discussions between the MI Department and the Beams Division management have begun to address operation of the Main Injector for Fixed Target running. We expect to start in a month on issues of extraction for SY120. For NuMI we are asked to provide 2.5E13/cycle from 5 Booster Batches. In addition to machine issues we need to document requirements for the beam permit (e.g. How to inhibit extraction of large emittance pulses?), lattices and other significant issues.
A problem has been observed with the time response of the MI Horiz IPM (non-magnetized). It is most apparent on cycles where protons are sent on round trip to/from the Recycler.
Martin Hu reported on RR status. We still see a tune change of 0.05 for $29 ramps. Problems with the flying wire are attributed to use of chromium plating on the carbon fiber in place of the intended gold plating. Brajesh reported that the the purchase orders for 70 DDC boards for the BPM system are moving through the purchasing process.